# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v2

@@ 2.3 Policy optimisation algorithms
RPD (Wang et al.,, 2022) primal-dual, projected descent with robust TD stationary point $\mathcal{O}(1/T^{1/4})$ /
RC-PO (Sun et al.,, 2024) trust region with projection general feasible improvement / /
Robust Sample-based PMD-PD (ours) see above softmax global RCMDP optimum / $\tilde{\mathcal{O}}(1/T^{1/3})$
Preliminaries
@@
The setting of the paper is based on robust constrained Markov Decision Processes (RCMDPs). Formally, an RCMDP is given by a tuple $(\mathcal{S},\mathcal{A},\rho,c_{0},c_{1:m},\mathcal{P},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\rho\in\Delta(\mathcal{S})$ is the starting distribution, $c_{0}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the cost function, $c_{i}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}\,,i\in[m]$ are constraint-cost functions, $\mathcal{P}$ is the uncertainty set, and $\gamma\in[0,1)$ is a discount factor. To denote the cardinality of sets $\mathcal{S}$ and $\mathcal{A}$, we use the convention $S:=|\mathcal{S}|$ and $A:=|\mathcal{A}|$. In RCMDPs, at any time step $l=0,\dots,\infty$, an agent observes a state $s_{l}\in\mathcal{S}$, takes an action $a_{l}\in\mathcal{A}$ based on its policy $\pi\in\Pi=(\Delta(\mathcal{A}))^{\mathcal{S}}$, and receives cost $c_{0}(s_{l},a_{l})$ and constraint-cost signals $c_{j}(s_{l},a_{l})$ for all $j\in[m]=\{1,\dots,m\}$. Following the action, the next state is sampled from a transition dynamics model $p\in\mathcal{P}$ according to $s_{l+1}\sim p(\cdot|s_{l},a_{l})$. The transition dynamics model $p$ is chosen from the uncertainty set $\mathcal{P}$ to solve a minimax problem. In particular, denoting
@@
The objective in Eq. 3 does not represent a single plausible transition kernel for the CMDP since the different suprema are usually met by different transition kernels. In practice, a choice is often made to take the supremum of the traditional cost only (Russel et al.,, 2020). Since this ignores the constraint-costs, one may instead optimise a related problem, namely the robust Lagrangian RCMDP objective (Bossens,, 2024),
@@
which only requires a single simulator and will allow optimisation to focus on constraint-costs if they are violated and on the traditional cost otherwise. If the inner optimisation problem is solved as $p_{k}$ at iteration $k$, this in turn corresponds to an equivalent value function over a traditional MDP (Taleghan and Dietterich,, 2018; Bossens,, 2024) according to the Lagrangian value function,
$\mathbf{V}_{\pi,p_{k}}(s)=\mathbb{E}\left[\sum_{l=0}^{\infty}\gamma^{l}\left(c_{0}(s_{l},a_{l})+\sum_{j=1}^{m}\lambda_{k,j}c_{j}(s_{l},a_{l})\right)\Big|s_{0}=s,\pi,p_{k},\right]\,$, (5)
which has a single cost function defined by $\mathbf{c}_{k}(s,a)=c_{0}(s,a)+\sum_{j=1}^{m}\lambda_{k,j}c_{j}(s,a)$ and no constraint-costs. The multiplier $\lambda_{k}$ denotes the vector of dual variables at iteration $k$ and $\lambda_{k,j}$ denotes the multiplier for constraint $j\in[m]$ at iteration $k$. Occasionally, in proofs where the iteration is not needed, the iteration index will be omitted.
@@
PMD-PD for Constrained MDPs
We briefly review the Policy Mirror Descent-Primal Dual (PMD-PD) algorithm (Liu et al., 2021b, ) for constrained MDPs. PMD-PD uses natural policy gradient with parametrised softmax policies, which is equivalent to the mirror descent.
@@
Analogous to 8, the regularised state-action value function is given by
@@
$F_{\lambda}:=1+\sum_{j=1}^{m}\lambda_{j}+\frac{\eta_{\lambda}m}{(1-\gamma)}\,$, (11)
@@
In this way, PMD-PD applies entropy regularisation with respect to the previous policy as opposed to the uniformly randomized policy used in Cen et al., (2022), allowing PMD-PD to converge to the optimal unregularised policy as opposed to the optimal regularised (and sub-optimal unregularised) policy.
@@
where the last step defines $\eta=\eta^{\prime}(1-\gamma)/d_{\rho}^{\pi_{k},p}(s)$ and subsequently removes the constant $\frac{1}{1-\gamma}d_{\rho}^{\pi_{k}^{t},p}(s)$ from the minimisation problem. Treating $\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}$ as any other value function, Eq. 4 is shown to be equivalent to the natural policy gradient update, as discussed by works in traditional MDPs (Zhan et al.,, 2023; Cen et al.,, 2022). In particular, based on Lemma 6 in Cen et al., (2022), the update rule derives from the properties of the softmax parametrisation in Eq. 15,
@@
Denoting $T=\sum_{k=1}^{K}t_{k}$, the oracle version (with perfect gradient info) converges at rate $\mathcal{O}(\log(T)/T)$ in value and constraint violation. PMP-PD Zero (oracle version aiming for zero constraint violation) achieves 0 violation with the same convergence rate for the value. Finally, a version with imperfect gradient information, the so-called sample-based version, was shown to converge at a rate $\tilde{\mathcal{O}}(1/T^{1/3})$, where $T$ denotes the total number of samples, thereby improving on model-free results which have a convergence rate $\mathcal{O}(1/T^{1/4})$.
Robust Sample-based PMD-PD for Robust Constrained MDPs
Having reviewed the performance guarantees of PMD-PD, we now derive the Robust Sample-based PMD-PD algorithm for robust constrained MDPs. The algorithm performs robust training based on Transition Mirror Ascent (TMA) (Wang et al.,, 2024) over the robust Lagrangian objective (Bossens,, 2024). Using TMA provides a principled way to compute adversarial dynamics while maintaining the transition dynamics within the bounds of the uncertainty sets. Such uncertainty sets allow for rich parametrisation such as Gaussian mixtures, which can estimate rich probability distributions, and entropy-based parametrisations, which are motivated by the optimal form of KL-divergence based uncertainty sets (see Table 7 for these examples). The algorithm takes place in the sample-based setting, where the state(-action) values are approximated from a limited number of finite trajectories. With a regret bound $\tilde{\mathcal{O}}(1/T^{1/3})$, results are in line with the Sample-based PMD-PD algorithm. The resulting algorithm is given in Algorithm 2. While the first subsections develop theory for the oracle setting, where exact values and policy gradients are given, the subsequent subsections derive convergence results for the sample-based setting, where values and policy gradients are estimates. In particular, Section 5.3 proposes and analyses an algorithm for approximate TMA and Section 5.4) proposes and analyses Robust Sample-based PMD-PD. The last subsection then proposes a practical implementation.
@@ 5.1 General assumptions
We assume that the dual variables are bounded by a constant, such that also $F_{\lambda}=\mathcal{O}(1)$, which allows bounding the regret of a policy with respect to the optimal Lagrangian in Eq. 4 to be bounded. The variables can take a wider range compared to other primal-dual algorithms due to the use of the augmented Lagrangian instead of clipping. The parameter is expected to be large for domains with highly active, restrictive constraints. The assumption is not particularly strong since it may be that $B_{\lambda}\gg\frac{2}{(1-\gamma)\zeta}$, the clipping bound usually taken for primal-dual variables (Ding et al.,, 2020). While with the modified Lagrangian update, the iterates obtained in the algorithm are always bounder, Assumption 2 additionally allows bounding the maximising dual variables of each iterate. In the sample complexity analysis of Robust Sample-based PMD-PD, the assumption is only required to limit the average regret induced by a mismatch of the dual variable with the optimal ones for that macro-iteration (see Lemma 28. There, the assumption could be removed when the dual variables are, on average, close to their optima of their respective inner problems.
@@
This assumption is common in MDPs. It ensures that when sampling from some $\mu$ instead of the initial state distribution $\rho$, any importance sampling corrections with $\mu$ and $d_{\mu}^{\pi}$ are finite. In particular, it ensures that for any transition kernel $p$ and any policy $\pi$ that the mismatch coefficient $M_{p}(\pi):=\left\lVert\frac{d_{\mu}^{\pi,p}}{\mu}\right\rVert_{\infty}$ is finite (Mei et al.,, 2020). This is also useful for robust MDPs (Wang et al.,, 2024) since
@@
is also finite. Similarly, via Lemma 9 of Mei et al., (2020), it ensures $U_{p}=\inf_{s\in\mathcal{S},k\geq 1}\pi_{k}(a_{p}^{*}(s)|s)>0$ when following softmax policy gradient, where $a_{p}^{*}(s)$ is the action for state $s$ under the optimal deterministic policy under the transition dynamics $p$. Again, this also finds its use in robust MDPs (Wang et al.,, 2024) where it follows that
@@
In traditional CMDPs, Slater’s condition implies strict feasibility of the constraints, zero duality gap, and complementary slackness for any convex optimisation problem, of which the CMDP is an instance due to its equivalence to a linear programming formulation (Altman,, 1999). The assumption of Slater points is common in CMDPs, and it is inherited from (Liu et al., 2021a, ). It is a ubiquitous but often hidden assumption when using dual and primal-dual algorithms, since it is needed to satisfy strong duality, such that $\min_{\pi\in\Pi}\max_{\lambda\geq 0}\mathbf{V}_{\pi,p}(\rho;\lambda)=\max_{\lambda\geq 0}\min_{\pi\in\Pi}\mathbf{V}_{\pi,p}(\rho;\lambda)$. Many constrained RL algorithms do mention the assumption explicitly (Achiam et al.,, 2017; Liu et al., 2021b, ; Ding et al.,, 2020, 2021). Note that we do not require knowledge of the parameter $\zeta$ while other primal-dual algorithms (Ding et al.,, 2020, 2021) need to know the parameter to clip the Lagrangian. Another approach, which assumes a Slater point but also does not require knowledge of $\zeta$ (Germano et al.,, 2024) bounds the dual variables by showing a strong no-regret property for both the primal and the dual regret minimizers. Approaches without Slater’s condition are possible using primal linear programming (Altman,, 1999; Efroni et al.,, 2020). Unfortunately, such techniques are less practical due to requiring estimates of the occupancy measure.
Assumption 4 is slightly more restrictive than in traditional CMDPs due to the requirement of the condition to hold for all $p\in\mathcal{P}$. The condition of having a subset of the policies being able to strictly satisfy all the constraints for all the transition dynamics reflects the RCMDP objective in Eq. 3 since it is presumed that the worst-case dynamic with respect to the constraints is solvable. For our purposes, the assumption ensures each of the individual CMDPs in the uncertainty set can be individually solved without duality gap similar to Liu et al., 2021b and that the dual variable optimum and iterates (see Lemma 25 and Lemma 27) can be bounded by a constant that can be significantly smaller than the $B_{\lambda}$ that bounds the space of dual variables.
Due to the occupancy being a function of an induced transition kernel which is potentially different for each policy, a zero duality gap may be hard to guarantee for RCMDPs even with Slater points (Wang et al.,, 2022). However, our analysis derives regret bounds with respect to the (primal) global RCMDP optimum. Our sample-based analysis provides a regret upper bound with respect to the (primal) global optimum of the RCMDP $(\pi^{*},p^{*},\lambda^{*})$ using Lemma 8, which is derived based on information from the iteration of the algorithm, namely the error tolerance of LTMA, the suboptimality of the dual variables, and the Lagrangian at $(\pi^{*},p_{k},\lambda_{k-1})$. Such errors are then shown to vanish in an average regret analysis.
@@
The convex set assumption ensures updates to the transition kernel will remain in the interior so long as it is on a line between two interior points, a common assumption for mirror descent which is also needed for TMA (Wang et al.,, 2024). The assumption is a standard (often hidden) assumption in gradient descent/ascent type algorithms to ensure the parameter remains in the feasible set.
The rectangular uncertainty set assumption is satisfied for most common uncertainty sets, e.g. $\ell_{1}$ and $\ell_{2}$ sets centred around a nominal value. The use of rectangularity assumptions allows us to efficiently use performance difference lemmas for transition kernels as well as make use of results for the optimality of TMA. The rectangularity assumption is at the basis of the proofs of our sample-based analysis, which uses it to bound the error based on an uncertainty budget for an $\ell_{q}$ uncertainty set. While Wang et al., (2024) mentions the policy optimisation part of their algorithm works with non-rectangular uncertainty sets in principle, the TMA algorithm only applies to rectangular uncertainty sets. Proposals for non-rectangular uncertainty sets do exist but the results are generally not so encouraging. Wiesemann et al., (2013) shows that solving the inner loop of both non-rectangular and non-convex uncertainty sets are both NP hard problems. Recent work demonstrates an actor-critic algorithm with an $\mathcal{O}(\epsilon^{-4})$ bound for non-rectangular uncertainty sets (Li et al.,, 2023). Other alternatives include $r$-rectangular uncertainty sets, which could help reduce the number of distinct factors included to compose the uncertainty set. However, recent work shows that $r$-rectangular and non-rectangular frameworks are only tractable if one assumes costs with no next-state dependencies, an assumption which undermines their usefulness (Grand-Clément et al.,, 2024).
Note that when using such parametrised transition kernels, it is convenient to formulate an uncertainty set in the parameter space, which we denote as $\mathcal{U}_{\xi}$. A few examples of parametrised transition kernels (PTKs) and their associated parametrised uncertainty sets can be found in Table 7, including the Entropy PTK (Wang et al.,, 2024), which presents a parametrisation related to the optimal parametrisation for KL divergence uncertainty sets of the form $\mathcal{P}_{s,a}=\{p:D_{\text{KL}}(p,\bar{p}(\cdot|s,a))\leq\kappa\}$ (Nilim and Ghaoui,, 2005), and the Gaussian mixture PTK (Wang et al.,, 2024), which is able to capture a rich set of distributions for each state-action pair.

### v4

@@ 2.3 Policy optimisation algorithms
(Wang et al., 2022) primal-dual, projected descent with robust TD rRC general stationary point $\mathcal{O}(1/T^{1/4})$ /
(Sun et al., 2024) trust region with projection rRC general feasible improvement / /
Ours double-loop primal-dual mirror descent/ascent rRC softmax global optimum / $\tilde{\mathcal{O}}(1/T^{1/3})$
Ours double-loop primal-dual mirror descent/ascent nRC softmax global with error $\delta_{\mathcal{P}}$ / $\tilde{\mathcal{O}}(1/T^{1/5})$
3 Preliminaries
@@
The setting of the paper is based on RCMDPs. Formally, an RCMDP is given by a tuple $(\mathcal{S},\mathcal{A},\rho,c_{0},c_{1:m},\mathcal{P},\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\rho\in\Delta(\mathcal{S})$ is the starting distribution, $c_{0}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ is the cost function, $c_{i}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}\,,i\in[m]$ are constraint-cost functions, $\mathcal{P}$ is the uncertainty set, and $\gamma\in[0,1)$ is a discount factor. To denote the cardinality of sets $\mathcal{S}$ and $\mathcal{A}$, we use the convention $S:=|\mathcal{S}|$ and $A:=|\mathcal{A}|$. In RCMDPs, at any time step $l=0,\dots,\infty$, an agent observes a state $s_{l}\in\mathcal{S}$, takes an action $a_{l}\in\mathcal{A}$ based on its policy $\pi\in\Pi\subseteq(\Delta(\mathcal{A}))^{\mathcal{S}}$, and receives cost $c_{0}(s_{l},a_{l})$ and constraint-cost signals $c_{j}(s_{l},a_{l})$ for all $j\in[m]=\{1,\dots,m\}$. Following the action, the next state is sampled from a transition dynamics model $p\in\mathcal{P}$ according to $s_{l+1}\sim p(\cdot|s_{l},a_{l})$. The transition dynamics model $p$ is chosen from the uncertainty set $\mathcal{P}$ to solve a minimax problem. In particular, denoting
@@
The objective in Eq. 3 does not represent a single plausible transition kernel for the CMDP since the different suprema are usually met by different transition kernels. In practice, a choice is often made to take the supremum of the traditional cost only (Russel et al., 2020). Since this ignores the constraint-costs, one may instead optimise a related problem, namely the robust Lagrangian RCMDP objective (Bossens, 2024),
@@
which only requires a single simulator and will allow optimisation to focus on constraint-costs if they are violated and on the traditional cost otherwise. If the inner optimisation problem is solved as $p_{k}$ at iteration $k$, this in turn corresponds to an equivalent value function over a traditional MDP (Taleghan & Dietterich, 2018; Bossens, 2024) according to the Lagrangian value function,
$\mathbf{V}_{\pi,p_{k}}(\rho)=\mathbb{E}\left[\sum_{l=0}^{\infty}\gamma^{l}\left(c_{0}(s_{l},a_{l})+\sum_{j=1}^{m}\lambda_{k,j}c_{j}(s_{l},a_{l})\right)\Big|s_{0}\sim\rho,\pi,p_{k},\right]\,$, (5)
which has a single cost function defined by $\mathbf{c}_{k}(s,a)=c_{0}(s,a)+\sum_{j=1}^{m}\lambda_{k,j}c_{j}(s,a)$ and no constraint-costs. The multiplier $\lambda_{k}$ denotes the vector of dual variables at iteration $k$ and $\lambda_{k,j}$ denotes the multiplier for constraint $j\in[m]$ at iteration $k$. Occasionally, in proofs where the iteration is not needed, the iteration index will be omitted. In the context of dynamic updates to $\lambda$, we often emphasise the value of $\lambda$ by including it as a pararameter in the notation $\mathbf{V}_{\pi,p_{k}}(\rho;\lambda)$.
@@
For any optimisation problem $\min_{x\in X}f(x)$, we define three notions of regret:
$f(x_{k})-f(x^{*})$ (regret)
$\sum_{k=1}^{K}f(x_{k})-f(x^{*})$ (cumulative regret)
$\frac{1}{K}\sum_{k=1}^{K}f(x_{k})-f(x^{*})$ (average regret)
where $\{x_{k}\}_{k\geq 1}$ represent the iterates of the algorithm and $x^{*}$ is the optimum. Based on these regret notions, the paper provides a sample-based analysis in terms of how many samples are required until the regret is below some small threshold $\epsilon>0$.
A first objective of interest is the maximisation over transition kernels within Eq. 3, i.e.
$\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi,p}(\rho;\lambda)\,$, (8)
where $\lambda\in\mathbb{R}^{m}$ and $\pi\in\Pi$ are fixed. Our analysis aims to show that it is possible to obtain, within a limited number of samples using only approximate evaluations of the Lagrangian w.r.t. the current parameters, a near-optimal maximiser, i.e. a transition kernel with a limited regret,
$\sup_{p\in\mathcal{P}}\mathbf{V}_{\pi,p}(\rho;\lambda)-\mathbf{V}_{\pi,p^{t}}(\rho;\lambda)\leq\epsilon\,$, (9)
where $p^{t}$ is the transition kernel being optimised at iteration $t\geq 1$.
A second objective of interest is the minimax optimisation problem in Eq. 3 itself. To design an optimal algorithm for the full RCMDP problem, the main goal of the sample-based analysis is to minimise the average regret w.r.t. Eq. 4,
$\frac{1}{K}\sum_{k=1}^{K}\Phi(\pi_{k})-\Phi(\pi^{*})\leq\epsilon\,$, (10)
within a limited number of samples. To count the samples, we account for those used for the evaluation and optimisation of the policy and the transition kernel, for a full sample-based analysis.
4 PMD-PD for CMDPs
We briefly review the PMD-PD algorithm for CMDPs (Liu et al., 2021b). PMD-PD uses natural policy gradient with parametrised softmax policies, which is equivalent to the mirror descent.
@@ 4 PMD-PD for CMDPs
Analogous to 11, the regularised state-action value function is given by
@@
$F_{\lambda}:=1+\sum_{j=1}^{m}\lambda_{j}+\frac{\eta_{\lambda}m}{(1-\gamma)}$ (14)
@@
In this way, PMD-PD applies entropy regularisation with respect to the previous policy as opposed to the uniformly randomized policy used in Cen et al. (2022), allowing PMD-PD to converge to the optimal unregularised policy as opposed to the optimal regularised (and sub-optimal unregularised) policy.
@@
where the last step defines $\eta=\eta^{\prime}(1-\gamma)/d_{\rho}^{\pi_{k},p}(s)$ and subsequently removes the constant $\frac{1}{1-\gamma}d_{\rho}^{\pi_{k}^{t},p}(s)$ from the minimisation problem. Treating $\tilde{\mathbf{Q}}_{\pi_{k}^{t},p}$ as any other value function, Eq. 4 is shown to be equivalent to the natural policy gradient update, as discussed by works in traditional MDPs (Zhan et al., 2023; Cen et al., 2022). In particular, based on Lemma 6 in Cen et al. (2022), the update rule derives from the properties of the softmax parametrisation in Eq. 18,
@@
Denoting $T=\sum_{k=1}^{K}t_{k}$, the oracle version (with perfect gradient info) converges at rate $\mathcal{O}(\log(T)/T)$ in value and constraint violation. PMP-PD Zero (oracle version aiming for zero constraint violation) achieves 0 violation with the same convergence rate for the value. Finally, the sample-based version, which has imperfect gradient information, was shown to converge at a rate $\tilde{\mathcal{O}}(1/T^{1/3})$, where $T$ denotes the total number of samples.
5 Robust Sample-based PMD-PD for RCMDPs
Having reviewed the performance guarantees of PMD-PD, we now derive the Robust Sample-based PMD-PD algorithm for RCMDPs. The algorithm performs robust training based on Transition Mirror Ascent (TMA) (Wang et al., 2024) over the robust Lagrangian objective (Bossens, 2024). Using TMA provides a principled way to compute adversarial dynamics while maintaining the transition dynamics within the bounds of the uncertainty sets. Such uncertainty sets allow for rich parametrisation such as Gaussian mixtures, which can estimate rich probability distributions, and entropy-based parametrisations, which are motivated by the optimal form of KL-divergence based uncertainty sets (see Table 7 for these examples). The algorithm takes place in the sample-based setting, where the state(-action) values are approximated from a limited number of finite trajectories. With an average regret bound $\tilde{\mathcal{O}}(1/T^{1/3})$, results are in line with the Sample-based PMD-PD algorithm. The resulting algorithm is given in Algorithm 3. While the first subsection develops the theory of TMA in the oracle setting (Section 5.2), where exact values and policy gradients are given, the subsequent subsections derive convergence results for the sample-based setting, where values and policy gradients are estimates. In particular, Section 5.3 proposes and analyses Approximate TMA, an algorithm for TMA within the sample-based setting, and Section 5.4) proposes and analyses Robust Sample-based PMD-PD in the full sample-based RCMDP setting. Further extensions are then considered, namely non-rectangular uncertainty sets (Section 5.5) and continuous state-action spaces (Section 5.6). The last subsection (Section 5.7) then proposes a practical implementation.
@@ 5.1 General assumptions
We assume that the dual variables are bounded by a constant, such that also $F_{\lambda}=\mathcal{O}(1)$, which allows bounding the regret of a policy with respect to the optimal Lagrangian in Eq. 4 to be bounded. The variables can take a wider range compared to other primal-dual algorithms due to the use of the augmented Lagrangian instead of clipping. The parameter is expected to be large for domains with highly active, restrictive constraints. The assumption is not particularly strong since it may be that $B_{\lambda}\gg\frac{2}{(1-\gamma)\zeta}$, the clipping bound usually taken for primal-dual variables (Ding et al., 2020). While with the modified Lagrangian update, the iterates obtained in the algorithm are always bounder, Assumption 2 additionally allows bounding the maximising dual variables of each iterate. In the sample complexity analysis of Robust Sample-based PMD-PD, the assumption is only required to limit the average regret induced by a mismatch of the dual variable with the optimal ones for that macro-iteration (see Lemma 29. There, the assumption could be removed when the dual variables are, on average, close to their optima of their respective inner problems.
@@
This assumption is common in MDPs. It ensures that when sampling from some $\mu$ instead of the initial state distribution $\rho$, any importance sampling corrections with $\mu$ and $d_{\mu}^{\pi}$ are finite. In particular, it ensures that for any transition kernel $p$ and any policy $\pi$ that the mismatch coefficient $M_{p}(\pi):=\left\lVert\frac{d_{\mu}^{\pi,p}}{\mu}\right\rVert_{\infty}$ is finite (Mei et al., 2020). This is also useful for RMDPs (Wang et al., 2024) since
@@
is also finite. Similarly, via Lemma 9 of Mei et al. (2020), it ensures $U_{p}=\inf_{s\in\mathcal{S},k\geq 1}\pi_{k}(a_{p}^{*}(s)|s)>0$ when following softmax policy gradient, where $a_{p}^{*}(s)$ is the action for state $s$ under the optimal deterministic policy under the transition dynamics $p$. Again, this also finds its use in RMDPs (Wang et al., 2024) where it follows that
@@
In traditional CMDPs, Slater’s condition implies strict feasibility of the constraints, zero duality gap, and complementary slackness for any convex optimisation problem, of which the CMDP is an instance due to its equivalence to a linear programming formulation (Altman, 1999). The assumption of Slater points is common in CMDPs, and it is inherited from (Liu et al., 2021a). It is a ubiquitous but often hidden assumption when using dual and primal-dual algorithms, since it is needed to satisfy strong duality, such that $\min_{\pi\in\Pi}\max_{\lambda\geq 0}\mathbf{V}_{\pi,p}(\rho;\lambda)=\max_{\lambda\geq 0}\min_{\pi\in\Pi}\mathbf{V}_{\pi,p}(\rho;\lambda)$. Many constrained RL algorithms do mention the assumption explicitly (Achiam et al., 2017; Liu et al., 2021b; Ding et al., 2020; 2021). Note that we do not require knowledge of the parameter $\zeta$ while other primal-dual algorithms (Ding et al., 2020; 2021) need to know the parameter to clip the Lagrangian. Another approach, which assumes a Slater point but also does not require knowledge of $\zeta$ (Germano et al., 2024) bounds the dual variables by showing a strong no-regret property for both the primal and the dual regret minimizers. Approaches without Slater’s condition are possible using primal linear programming (Altman, 1999; Efroni et al., 2020). Unfortunately, such techniques are less practical due to requiring estimates of the occupancy measure.
Assumption 4 is slightly more restrictive than in traditional CMDPs due to the requirement of the condition to hold for all $p\in\mathcal{P}$. The condition of having a subset of the policies being able to strictly satisfy all the constraints for all the transition dynamics reflects the RCMDP objective in Eq. 3 since it is presumed that the worst-case dynamic with respect to the constraints is solvable. For our purposes, the assumption ensures each of the individual CMDPs in the uncertainty set can be individually solved without duality gap similar to Liu et al. (2021b) and that the dual variable optimum and iterates (see Lemma 26 and Lemma 28) can be bounded by a constant that can be significantly smaller than the $B_{\lambda}$ that bounds the space of dual variables.
Due to the occupancy being a function of an induced transition kernel which is potentially different for each policy, a zero duality gap may be hard to guarantee for RCMDPs even with Slater points (Wang et al., 2022). However, our sample-based analysis provides an average regret upper bound with respect to the (primal) global optimum of the RCMDP $(\pi^{*},p^{*},\lambda^{*})$ using Lemma 8, which is derived based on information from the iteration of the algorithm, namely the error tolerance of LTMA, the suboptimality of the dual variables, and the Lagrangian at $(\pi^{*},p_{k},\lambda_{k-1})$. Such errors are then shown to vanish in an average regret analysis.
@@
The convex set assumption ensures updates to the transition kernel will remain in the interior so long as it is on a line between two interior points, a common assumption for mirror descent which is also needed for TMA (Wang et al., 2024). The assumption is a standard (often hidden) assumption in gradient descent/ascent type algorithms to ensure the parameter remains in the feasible set.
The rectangular uncertainty set assumption is satisfied for most common uncertainty sets, e.g. state-action conditioned $\ell_{1}$ and $\ell_{2}$ sets centred around nominal probability distributions. In our main theorem, a tight rectangular $\ell_{q}$ uncertainty set is used in Lemma 24 to provide an improved upper bound on the average regret, although the same convergence rate is achieved for general uncertainty sets albeit with a larger constant. While Wang et al. (2024) mentions the policy optimisation part of their algorithm works with non-rectangular uncertainty sets in principle, the guarantees of the TMA algorithm only apply to rectangular uncertainty sets, where it can provide an ascent property to guarantee an increase in the objective. Consequently, Lagrangian TMA algorithms inherit this property.
Proposals for solving the inner problem of non-rectangular uncertainty sets do exist but most results are not encouraging. As noted by Wiesemann et al. (2013), the inner loops of non-rectangular and non-convex uncertainty sets are NP-Hard problems. Other alternatives include $r$-rectangular uncertainty sets, which could help reduce the number of distinct factors included to compose the uncertainty set. Recent work shows that $r$-rectangular and non-rectangular frameworks are only tractable if one assumes costs with no next-state dependencies, an assumption that may undermine their usefulness (Grand-Clément et al., 2024).
However, recent research has also highlighted some benefits of non-rectangular uncertainty sets, such as allowing statistically optimal confidence regions (Li et al., 2023) and covering the maximum likelihood areas rather than the edges (Kumar et al., 2025). These works also introduce techniques for approximately solving the inner and outer problems in the context of RMDPs. To supplement our analysis of rectangular sets, we also consider non-rectangular RCMDPs in Section 5.5. The analysis uses the conservative policy iteration algorithm as defined for transition kernels in Li et al. (2023), which treats policy iteration as a special case of Approximate TMA.
Note that when using parametrised transition kernels, it is convenient to formulate an uncertainty set in the parameter space, which for a parameter $\xi$ we denote as $\mathcal{U}_{\xi}$. A few examples of parametrised transition kernels (PTKs) and their associated parametrised uncertainty sets can be found in Table 7 (see Appendix A), including the Entropy PTK (Wang et al., 2024), which presents a parametrisation related to the optimal parametrisation for KL divergence uncertainty sets of the form $\mathcal{P}_{s,a}=\{p:D_{\text{KL}}(p,\bar{p}(\cdot|s,a))\leq\kappa\}$ (Nilim & Ghaoui, 2005), and the Gaussian mixture PTK (Wang et al., 2024), which is able to capture a rich set of distributions for each state-action pair. Such uncertainty sets can be considered as general uncertainty sets although they are often also contained within an $\ell_{q}$ set.